Circular No. 6558

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997Z IN NGC 3261 R. Martin, A. Williams, and S. Woodings, Perth Observatory, report

the discovery of a supernova (mag R ∼ 15.5) from a CCD frame taken on Feb. 10.64 UT. SN 1997Z is located 58′′ east and 10′′ south of the center of NGC 3261 (α = 10^{h}29^{m}.0, δ = −44^{o}39′, equinox 2000.0).

S. Benetti and M. Turatto, European Southern Observatory (ESO); and E. Cappellaro, Osservatorio di Padova, communicate: “Inspection of a reduced CCD spectrogram (range 380–800 nm, resolution 1.1 nm) obtained on Feb. 12.12 UT with the ESO/MPI 2.2-m telescope (+ EFOSC2) confirms SN 1997Z as a type-II supernova, ∼ 100 days after explosion. In the spectrum, several lines with typical P-Cyg profiles are superimposed on a relatively red continuum. Most of the lines arise from H I, Na I, Fe II, and Sc II. The expansion velocities (in km/s) deduced from the minima of the strongest lines are 7200 (Hα), 6150 (Na I); 4800 (Hβ), and 3550 (Fe II). With the same instrument, we obtained the following magnitudes for the new object: B = 19.0, B − V = +1.4, and V − R = +1.3. The offsets from the galaxy nucleus are 55′′.1 east and 8′′.7 south.”

V1333 AQUILAE A. M. Levine, Massachusetts Institute of Technology (MIT), and B.

Thomas, Goddard Space Flight Center (GSFC), on behalf on the RXTE ASM team at GSFC and MIT; M. Garcia and P. Berlind, Smithsonian Astrophysical Observatory; and P. Callanan, University College Cork, report that the low-mass x-ray binary Aquila X-1 (= V1333 Aql) appears to be in a moderate-intensity outburst in quick-look results from the RXTE All Sky Monitor. Recent daily-average flux levels in the band 2–12 keV are as follows: Jan. 27, ∼ 60 mCrab; 28, 120; 29, 190; 30, 260; Feb. 1, 310; 2–5, ∼ 360. Observations of the optical counterpart V1333 Aql from the Whipple Observatory 1.2-m telescope on Feb. 6.543 UT yield V = 16.9. Observations at other wavelengths are encouraged.

COMET 46P/WIRTANEN Total visual magnitude estimates: Jan. 3.09 UT, 12.6 (A. Hale, Cloud-

croft, NM, 0.41-m reflector); Feb. 1.75, 11.8 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 4.09, 11.3 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 6.74, 11.4 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector).

1997 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green